OpenStudy (anonymous):

Use the properties of trig functions to find the exact value: sin(-pi/12)csc(25pi/12)?

OpenStudy (anonymous):

\[\sin{\frac{-\pi}{12}} \csc{\frac{25\pi}{12}} \\ -\sin{\frac{\pi}{12}} \csc{\frac{\pi}{12}} \\ -\sin{\frac{\pi}{12}} \frac{1}{\sin{\frac{\pi}{12}}} \\ -1 \]

OpenStudy (anonymous):

note that \[\csc{\frac{25\pi}{12}} = \csc{(2\pi+\frac{\pi}{12})}= \csc{(\frac{\pi}{12})}\]

OpenStudy (anonymous):

ahhh i see the csc\[25\pi \div12\] can have 24pi taken out so its just csc pi/12 and its the reciprocal of 1/sin pi/12, then cancel the two out and dont forget the negative sign, so its -1. got it. Thanks!

OpenStudy (anonymous):

idk why i forgot about the periodic property!

OpenStudy (anonymous):

>.> that happened to me when I do my exam..

OpenStudy (anonymous):

yeah its tough memorizing all this in a 6 week accelerated precalc class 0_0 lol